Sol Summary Report – February 27th

Crew 275 Sol Summary Report 27Feb2023

Sol: 15

Summary Title: Cosmic Chasma

Author’s name: Jeremy Rabineau, Commander.

Mission Status: Nominal.

Sol Activity Summary:

Today we performed our very first 4-hour EVA at Candor Chasma. The aim of this ambitious EVA was to define checkpoints and perform photogrammetry of the area using Anafi drones. This time, the area of interest was much wider and required more pictures (the computer is still working on building the 3D render of the terrain). The EVA crew was amazed by the landscapes that they had the opportunity to see, and the four remaining crewmembers are looking forward to seeing that for themselves later during the week.

With the clear sky of last night, our crew astronomer thought that the observations he had planned would have been performed, but he was disappointed to see that it was not the case. Unfortunately, it was also too windy to observe the Sun today…

At the GreenHab, it was time for an intense harvest. New seeds have also been planted. We are all looking forward to seeing them grow as we get closer to the end of our mission every day.

Look Ahead Plan:

After today’s EVA, some post-processing is needed to prepare the two next EVAs to Candor Chasma. The first one would take place on Wednesday (Sol 17) and the second one on Friday (Sol 19). The EVA crews would have to find specific checkpoints based on a 2D map or a 3D render. We are also planning to have a second geology EVA on Sol 20, to further test the MetMet instrument. We would like to take the opportunity of these different EVAs to change the batteries of the weather instruments, since they cannot last for more than two days, especially in cold weather.

The BBC is planned to arrive at the MDRS on Friday (Sol 19) between 8:30 and 9:00. They would follow us on our EVA to Candor Chasma, then film inside the station in the afternoon and leave at around 17:00.

Anomalies in work: None.

Weather: Sunny and windy, low -2°C/28°F, high 8°C/46°F.

Crew Physical Status: Optimal.

EVA: None.

Reports to be filed: Sol Summary, EVA Report, Operations Report, Journalist Report, GreenHab Report, Astronomy Report.

Support Requested: None.

Copyright © The Mars Society. All rights reserved. | Main Site